Money is borrowed at 14% simple interest. after one year, $1281.36 pays off the loan. how much was originally borrowed?
1
Expert's answer
2014-06-16T02:02:48-0400
Simple interest is determined by multiplying the interest rate by the principal by the number of periods.
Simple interest = P*I*N,

where P is the loan amount, I is the interest rate, N the duration. In our case we have SI=1282.36$, I=14% and N=1 year.
1282.36=P*0.14*1+P; 1282.36=1.14*P; P=1124.88$.

Answer. 1124.88$.